OpenStudy (anonymous):

Bethany opened a store credit card to purchase a TV for $589. She put the entire purchase on the credit card. Her APR is 29.99% and the minimum payment is 5%. Assuming she makes only minimum payments, what is the remaining balance on her card after 3 months? $504.99 $532.24 $545.90 $574.27

OpenStudy (anonymous):

29.99% = .2999 .2999/12 = .02499 (rounded to .025) add 1 = 1.025 589 x 1.025 = 603.73 minimum payment is 5% of 603.73 .05 x 603.73 = 30.19 per month For 3 months it is (30.19 x 3) = 90.57 603.73 - 90.57 = 513.16 This IS NOT the right answer, but I don't know what I am doing wrong. I am trying to go by this... http://www.ehow.com/how_5064522_calculate-card-monthly-payments-apr.html somebody see if they can do this correctly...PLEASE

jimthompson5910 (jim_thompson5910):

There are 2 ways to do this, but unfortunately, none of them get to any of the answers here (it gets close to one though) ======================================================================================= Method # 1: Make monthly payment, then calculate interest on remaining balance Monthly Payment: 589 * 0.05 = 29.45 1st month Make Monthly Payment: 589 - 29.45 = 559.55 Add on interest: 559.55 + 29.99% = 559.55 + (1/12)*0.2999*559.55 = 573.53 Remaining Balance: 573.53 2nd month Make Monthly Payment: 573.53 - 29.45 = 544.08 Add on interest: 544.08 + 29.99% = 544.08 + (1/12)*0.2999*544.08 = 557.68 Remaining Balance: 557.68 3rd month Make Monthly Payment: 557.68 - 29.45 = 528.23 Add on interest: 528.23 + 29.99% = 528.23 + (1/12)*0.2999*528.23 = 541.43 Remaining Balance: 541.43 ======================================================================================= Method # 2: Calculate interest, then make monthly payment Monthly Payment: 589 * 0.05 = 29.45 month # 1 Add on interest: 589 + 29.99% = 589 + (1/12)*0.2999*589 = 603.72 Make Monthly Payment: 603.72 - 29.45 = 574.27 Remaining Balance: 574.27 month # 2 Add on interest: 574.27 + 29.99% = 574.27 + (1/12)*0.2999*574.27 = 588.62 Make Monthly Payment: 588.62 - 29.45 = 559.17 Remaining Balance: 559.17 month # 3 Add on interest: 559.17 + 29.99% = 559.17 + (1/12)*0.2999*559.17 = 573.14 Make Monthly Payment: 573.14 - 29.45 = 543.69 Remaining Balance: 543.69 =======================================================================================

OpenStudy (anonymous):

So if you were to pick an answer yours would be?

jimthompson5910 (jim_thompson5910):

well I'm getting close to $545.90 I'm not sure why it's not landing exactly on it but since it's off by a few dollars (which is not much), I'd say the answer is $545.90

jimthompson5910 (jim_thompson5910):

did they give you any formulas to work with?

OpenStudy (anonymous):

They did but i cant seem to find it. So ill just go with 545.90. Thanks alot for your time and help!

jimthompson5910 (jim_thompson5910):

and I'm using a bunch of other tools and calculators and I'm getting the same thing pretty much for instance, this is one tool I used http://www.calculator.net/credit-card-calculator.html?balance=589&rate=29.99&minimum=29.45&payoffoption=0&fixedpaymentamount=35&x=98&y=15 and it gives the remaining balance on month #3 to be 543.70, which is one cent off of what I got

jimthompson5910 (jim_thompson5910):

let me know if that answer of $545.90 is the correct one or not

OpenStudy (anonymous):

I was wondering how this was supposed to be done....I guess I was just " way off ". Sorry about that Blake95.

jimthompson5910 (jim_thompson5910):

i could be off too...or there may be a typo? idk

jimthompson5910 (jim_thompson5910):

seems odd though how the calculator is giving something way different from the answer choices

OpenStudy (anonymous):

Ill let you guys know if thats the right answer! And no problem at all Kellie! I appreciate all the free help!

OpenStudy (anonymous):

It was right!

jimthompson5910 (jim_thompson5910):

ok that's great, just got to figure out if that was a typo or if I was using the wrong method/formula